PLEASE HELP!!
find the value or values of a in the
domain of f for which f(a) equals the given number.
f(x)=x^2+2x-2; f(a)= -1

Answer by stanbon(75887)

find the value or values of a in the
domain of f for which f(a) equals the given number.
f(x)=x^2+2x-2; f(a)= -1
------
a^2 + 2a -2 = -1
a^2 + 2a -1 = 0
-----
a = [-2 +- sqrt(4-4*1*-1)]/2
-----
a = [-2 +- sqrt(8)]/2
-----
a = [-2 +- 2sqrt(2)]/2
-----
a = -1 + sqrt(2) or a = -1 - sqrt(2)
--------------
Cheers,
Stan H.
-------------